# On generalized Beauville decompositions

### v1

@@ 4.1. The universal Picard stack
Let $\mathfrak{M}_{g}$ be the moduli stack of (not necessarily stable) nodal curves of genus $g$ and let $\mathfrak{C}_{g}\to\mathfrak{M}_{g}$ be the universal curve. We consider the universal Picard stack $\mathfrak{Pic}_{g}\to\mathfrak{M}_{g}$ parameterizing total degree 00 line bundles. The relative Picard stack $\mathfrak{Pic}^{\mathrm{rel}}_{g}$ is the quotient of $\mathfrak{Pic}_{g}$ by relative inertia $B{\mathbb{G}}_{m}$ over $\mathfrak{M}_{g}$; in other words, it is the rigidification of the universal Picard stack with respect to the automorphism ${\mathbb{G}}_{m}$.
@@
Since this class is invariant under twisting ${\mathcal{L}}$ by the line bundle from the base, it descends to
@@
This is because the universal sheaf can be interpreted as an admissible line bundle on the quasi-stable model of the family of nodal curves by [18]. It is necessary to allow unstable nodal curves and line bundles which are not of multi-degree 00 for such a map to exist. In particular we have a natural morphism
@@ 4.2. Universal double ramification cycle relations
[…]
where $j:\mathcal{M}_{g}\hookrightarrow\overline{\mathcal{M}}_{g}^{\leq 1}$ is an open embedding, $\iota:{\mathcal{M}}_{g-1,2}\to\overline{\mathcal{M}}_{g}^{\leq 1}$ is the morphism gluing the two marked points on the curve, and $J_{g-1,2}^{-1}$ is the relative Jacobian of degree $-1$ line bundles over $\mathcal{M}_{g-1,2}$. The morphism $\epsilon$ is given by $(\widetilde{C},L)\mapsto(C,\nu_{*}L)$ where $\nu:\widetilde{C}\to C$ is the normalization of a nodal curve with one self-node. The image of $\iota\circ\epsilon$ is the singular locus of $\pi$. The relative Jacobian $J_{g-1,2}^{-1}$ is isomorphic to $J_{g-1,2}$ under a translation
@@
for some class $\alpha\in\bigoplus_{w<2g-2}\mathrm{CH}^{g-1}_{(w)}(J_{g-1,2},{\mathbb{Q}})$. Here $w$ stands for the weight with respect to the “multiplication by $N$” map on $J_{g-1,2}$.
@@
We restrict Pixton’s formula for $\mathsf{P}_{g}^{g+1}$ to the relative Picard stack over ¯$\overline{\mathcal{M}}_{g}^{\leq 1}$; we thus focus on the substack $\mathfrak{M}_{g}^{\leq 1}\subset\mathfrak{M}_{g}$ where the quasi-stable curve stabilizes to a stable curve with at most one self node. Let $i:\mathfrak{Pic}_{\delta}^{\mathrm{rel}}\to\mathfrak{Pic}^{\mathrm{rel}}_{g}$ be the inclusion of the boundary strata where the underlying curve has at least one node. For an integer $d$, let $i^{d}:\mathfrak{Pic}^{\mathrm{rel}}_{\Gamma_{d}}\to\mathfrak{Pic}^{\mathrm{rel}}_{g}$ be the boundary stratum corresponding to a decorated prestable graph $\Gamma_{d}$ with two vertices of degree splitting $d,-d$, connected by two edges $e_{1}=(h_{1},h_{1}^{\prime})$ and $e_{2}=(h_{2},h_{2}^{\prime})$. Then Pixton’s formula has the form
@@
For our purpose, we further refine the relation $\mathsf{P}_{g}^{g+1}=0$ using the “multiplication by $N$” map
$[N]:\mathfrak{Pic}_{g}^{\mathrm{rel}}\to\mathfrak{Pic}_{g}^{\mathrm{rel}}.$
@@
$[N]^{*}\theta^{g+1}=N^{2g+2}\cdot\theta^{g+1}$
@@
$[N]^{*}\left(\theta^{a}i_{*}\big{(}(\psi_{h}+\psi_{h^{\prime}})^{b}\big{)}\right)=N^{2a}\cdot\theta^{a}i_{*}\big{(}(\psi_{h}+\psi_{h^{\prime}})\big{)}^{b}$
@@
$[N]^{*}\left(\theta^{k}i^{d}_{*}\big{(}(\psi_{h_{1}}+\psi_{h_{1}^{\prime}})^{\ell}(\psi_{h_{2}}+\psi_{h_{2}^{\prime}})^{m}\big{)}\right)=N^{2k}\cdot\theta^{k}i^{d/N}_{*}\big{(}(\psi_{h_{1}}+\psi_{h_{1}^{\prime}})^{\ell}(\psi_{h_{2}}+\psi_{h_{2}^{\prime}})^{m}\big{)}$
if $N$ divides $d$ and zero otherwise. Applying $[N]^{*}$ to the right-hand side of (32), the first two terms are clearly polynomials in $N$ and for the third term, we have
@@
which is a polynomial in $N$ because $f_{k,\ell,m}(d)$ are polynomials. Therefore we find an expression of $[N]^{*}\mathsf{P}^{g+1}_{g}=0$ which is a polynomial in $N$, so that each coefficient gives rise to a relation on $\mathfrak{Pic}^{\mathrm{rel}}_{g}$. We take the coefficient of $N^{2g+2}$ in the expression of $[N]^{*}\mathsf{P}_{g}^{g+1}=0$ and pull it back to $\overline{J}_{g}^{\leq 1}$. In particular, the second summand of the right-hand side of (32) does not contribute for weight reasons. This gives the desired relation on $\overline{J}_{g}^{\leq 1}$ of the proposition.
@@ 1. $\bullet$
    genus $g-1$ with degree $-1$, and
@@ 2. $\bullet$
    genus 00 with degree $1$
connected by two edges corresponds to a class under the pushforward $\iota_{*}\epsilon_{*}t_{*}$. We denote the two edges by $e_{1}=(h_{1},h_{1}^{\prime})$ and $e_{2}=(h_{2},h_{2}^{\prime})$ where $h_{1},h_{2}$ are the two half-edges attached to the genus $g-1$ vertex. After pulling back to $\overline{J}_{g}^{\leq 1}$, other strata do not contribute by the quasi-stable model of a rank $1$ torsion-free sheaf on a nodal curve. Then we have
@@
We explain how to interpret the classes given by prestable graphs decorated by $\psi$-monomials as classes pushed forward from $J_{g-1,2}$. Let ${\mathcal{L}}$ be the universal line bundle on the universal curve over $J_{g-1,2}$ trivialized along the first marking, and let $\xi_{2}\in\mathrm{CH}^{1}_{(1)}(J_{g-1,2},{\mathbb{Q}})$ be $c_{1}({\mathcal{L}})$ pulled back along the second marking. Analogous classes are defined on $J_{g-1,2}^{-1}$. We take the coefficient of $N^{1}$ in the expression of $[N]^{*}\mathsf{P}^{1}_{0,(1,-1)}=0$. This gives a relation
@@
at the unstable genus 00 vertex. Since $\xi_{h^{\prime}_{1}}=\xi_{h_{1}}=0$ and $\xi_{h^{\prime}_{2}}=\xi_{h_{2}}=\xi_{2}$, the “other terms” in (33) have an expression in terms of monomials of $\theta,\psi_{\bullet},\xi_{2}$. Under the translation, these monomials pull back to classes on $J_{g-1,2}$ of weights $<2g-2$. ∎
@@
Although the “multiplication by $N$” map extends to the relative Picard stack $\mathfrak{Pic}^{\mathrm{rel}}_{g}$, the cohomology $H^{*}(\mathfrak{Pic}^{\mathrm{rel}}_{g},{\mathbb{Q}})$ may not be a direct sum of weight spaces since it is infinite-dimensional.
We denote by $J^{\underline{0}}_{g}\subset\overline{J}^{\leq 1}_{g}$ the relative Jacobian of degree 00 line bundles. We have the tautological relation
$\theta^{g+1}=0\in\mathrm{CH}^{g+1}(J^{\underline{0}}_{g},{\mathbb{Q}})$
@@
By Proposition 4.2, it is enough to calculate the pushforward along $\pi$ of the classes $\iota_{*}\epsilon_{*}t_{*}(\theta^{g-1}/(g-1)!)$ and $\iota_{*}\epsilon_{*}t_{*}\alpha$. Since $\alpha\in\mathrm{CH}^{g-1}(J_{g-1,2},{\mathbb{Q}})$ has weights $<2g-2$, we have $\pi_{*}\iota_{*}\epsilon_{*}t_{*}\alpha=\iota_{*}\pi^{\prime}_{*}t_{*}\alpha=0$. We also recall
@@
where $0:{\mathcal{M}}_{g-1,2}\to J_{g-1,2}$ is the 00-section; see e.g. [16]. Therefore by the diagram (31) we have
@@
We recall a few standard tautological classes. The cotangent line of the $i$-th marking defines a line bundle ${\mathbb{L}}_{i}$ on ¯$\overline{{\mathcal{M}}}_{g,n}$ with first Chern class
@@
After restriction, we have the $\kappa$-classes on ¯$\overline{\mathcal{M}}_{g}^{\leq 1}$. We also define the boundary class
@@
We consider $J^{\underline{0}}_{g}\subset\overline{J}_{g}^{\leq 1}$ the open locus of degree 00 line bundles. Since its complement has (complex) codimension $2$, the restriction $H^{2}(\overline{J}_{g}^{\leq 1},{\mathbb{Q}})\to H^{2}(J^{\underline{0}}_{g},{\mathbb{Q}})$ is an isomorphism.
We apply the long exact sequence
$\cdots\to H^{0}(J^{\underline{0}}_{g}\setminus J_{g},{\mathbb{Q}})\to H^{2}(J^{\underline{0}}_{g},{\mathbb{Q}})\to H^{2}(J_{g},{\mathbb{Q}})\to H^{1}(J^{\underline{0}}_{g}\setminus J_{g},{\mathbb{Q}})\to\cdots.$
By [17], $H^{2}(J_{g},{\mathbb{Q}})$ is spanned by $\theta$ and $\kappa_{1}$. The lemma follows since the image of $H^{0}(J^{\underline{0}}_{g}\setminus J_{g},{\mathbb{Q}})$ is spanned by $\delta$. ∎
@@
Let $\bar{p}:\overline{{\mathcal{C}}}_{g}\to J_{g-1,2}^{-1}$ be the universal curve on $\overline{J}_{g}^{\leq 1}$ pulled back to $J^{-1}_{g-1,2}$ along $\iota\circ\epsilon$, and let $q:\overline{{\mathcal{C}}}^{\mathrm{qs}}_{g}\to\overline{{\mathcal{C}}}_{g}$ be the quasi-stable model. Let ${\mathcal{L}}^{\mathrm{qs}}$ be the universal admissible line bundle on $\overline{{\mathcal{C}}}^{\mathrm{qs}}_{g}$ trivialized along the first marking. The normalization of $\overline{{\mathcal{C}}}^{\mathrm{qs}}_{g}$ is the disjoint union of a ${\mathbb{P}}^{1}$-bundle ${\mathbb{P}}$ over $J_{g-1,2}$ and the universal curve ${\mathcal{C}}_{g-1,2}$. By [18, eq. (8)], we have a short exact sequence
@@
where we used the self-intersection formula $[x_{i}]^{2}=x_{i*}(-\psi_{i})$. ∎
@@
By taking the weight $2g$ part with respect to the “multiplication by $N$” map, we get
@@
This shows that $a\kappa_{1}$ is supported on the boundary ¯$\overline{{\mathcal{M}}}_{g}^{\leq 1}\setminus{\mathcal{M}}_{g}$ and is therefore proportional to $\delta$ by the irreducibility of the boundary. [^8] If $g\geq 3$, we can directly deduce $a=0$ from $\kappa_{1}\neq 0\in H^{2}({\mathcal{M}}_{g},{\mathbb{Q}})$.
@@
Since the excess intersection term for $\iota$ is $-(\psi_{1}+\psi_{2})$, we have
@@
By taking the weight $22g-$ 2 part, we get
@@
Here the coefficient $-\frac{1}{48}$ can be uniquely determined by the relation
@@
Therefore, we have to consider the larger open subset ¯$\overline{{\mathcal{M}}}^{\mathrm{int}}_{2}\subset\overline{{\mathcal{M}}}_{2}$ consisting of integral curves and the relative compactified Jacobian $\pi^{\mathrm{int}}:\overline{J}^{\mathrm{int}}_{2}\to\overline{{\mathcal{M}}}^{\mathrm{int}}_{2}$ as in Theorem 4.1(b).
Since the codimension of ¯$\overline{\mathcal{M}}_{2}^{\mathrm{int}}\setminus\overline{{\mathcal{M}}}_{2}^{\leq 1}$ is $2$, Proposition 4.8(a) shows that a generalized theta divisor has to be of the form
@@
¯$\overline{{\mathcal{M}}}_{1,2}^{\circ}=\overline{{\mathcal{M}}}_{1,2}\setminus\left(\overline{{\mathcal{M}}}_{1,1}\times\overline{{\mathcal{M}}}_{0,3}\right)\subset\overline{{\mathcal{M}}}_{1,2}$
@@
On $\overline{J}^{\mathrm{int}}_{2}$, the universal double ramification cycle relation $\mathsf{P}^{3}_{2}=0$ contains contribution from the boundary strata with two loops. However, those boundary strata contribute to lower degree coefficients of the expression $[N]^{*}\mathsf{P}^{3}_{2}=0$. Therefore Proposition 4.2 continues to hold on $\overline{J}^{\mathrm{int}}_{2}$. A direct computation yields $\alpha_{(0)}=\frac{1}{480}(\psi_{1}+\psi_{2})$. Combining with Proposition 4.2 and Lemma 4.7, we obtain
00 0 $=\pi^{\mathrm{int}}_{*}(\theta{\Theta}^{3})=\pi^{\mathrm{int}}_{*}\left(\iota^{\mathrm{int}}_{*}\epsilon_{*}t_{*}\Big{(}\frac{1}{8}(\theta+\frac{1}{2}(\psi_{1}+\psi_{2}))^{2}+6\theta\alpha_{(0)}-\frac{3b}{8}\theta(\psi_{1}+\psi_{2})\Big{)}+3b^{2}\theta^{2}\delta^{2}\right)$

### v2

@@ 4.1. The universal Picard stack
Let $\mathfrak{M}_{g}$ be the moduli stack of (not necessarily stable) nodal curves of genus $g$ and let $\mathfrak{C}_{g}\to\mathfrak{M}_{g}$ be the universal curve. We consider the universal Picard stack $\mathfrak{Pic}_{g}\to\mathfrak{M}_{g}$ parameterizing total degree 0 line bundles. The relative Picard stack $\mathfrak{Pic}^{\mathrm{rel}}_{g}$ is the quotient of $\mathfrak{Pic}_{g}$ by relative inertia $B{\mathbb{G}}_{m}$ over $\mathfrak{M}_{g}$; in other words, it is the rigidification of the universal Picard stack with respect to the automorphism ${\mathbb{G}}_{m}$.
@@
Since this class is invariant under twisting $\mathcal{L}$} by a line bundle from the base, it descends to
@@
This is because the universal sheaf on $C\times_{B}\overline{J}_{C}$ can be interpreted as an admissible line bundle on the quasi-stable model of the family of nodal curves by [19]. It is necessary to allow unstable nodal curves and line bundles which are not of multi-degree 0 for such a map to exist. In particular we have a natural morphism
@@ 4.2. Universal double ramification cycle relations
(32) $J_{g-1,2}$$J_{g-1,2}^{-1}$$\overline{J}_{\delta}$$\overline{J}_{g}^{\leq 1}$$J_{g}$$\mathcal{M}_{g-1,2}$$\overline{\mathcal{M}}_{g}^{\leq 1}$$\mathcal{M}_{g}$$\scriptstyle{t}$$\scriptstyle{\epsilon}$$\scriptstyle{\pi^{\prime}}$$\scriptstyle{\iota}$$\scriptstyle{\pi}$$\scriptstyle{j}$$\scriptstyle{\iota}$$\scriptstyle{j}$
where $j:\mathcal{M}_{g}\hookrightarrow\overline{\mathcal{M}}_{g}^{\leq 1}$ is the natural open embedding, $\iota:{\mathcal{M}}_{g-1,2}\to\overline{\mathcal{M}}_{g}^{\leq 1}$ is the morphism gluing the two marked points on the curve, and $J_{g-1,2}^{-1}$ is the relative Jacobian of degree 1-1 line bundles over $\mathcal{M}_{g-1,2}$. The morphism $\epsilon$ is given by $(\widetilde{C},L)\mapsto(C,\nu_{*}L)$ where $\nu:\widetilde{C}\to C$ is the normalization of a nodal curve with one self-node. The image of $\iota\circ\epsilon$ is the singular locus of $\pi$. The relative Jacobian $J_{g-1,2}^{-1}$ is isomorphic to $J_{g-1,2}$ under a translation
@@
for some class $\alpha\in\bigoplus_{w<2g-2}\mathrm{CH}^{g-1}_{(w)}(J_{g-1,2},{\mathbb{Q}})$. Here $w$ stands for the weight with respect to the “multiplication by NN” map on $J_{g-1,2}$.
@@
We restrict Pixton’s formula for $\mathsf{P}_{g}^{g+1}$ to the relative Picard stack over $\overline{\mathcal{M}}_{g}^{\leq 1}$; we thus focus on the substack $\mathfrak{M}_{g}^{\leq 1}\subset\mathfrak{M}_{g}$ where the quasi-stable curve stabilizes to a stable curve with at most one self-node. Let $i:\mathfrak{Pic}_{\delta}^{\mathrm{rel}}\to\mathfrak{Pic}^{\mathrm{rel}}_{g}$ be the inclusion of the boundary strata where the underlying curve has at least one node. For an integer $d$, let $i^{d}:\mathfrak{Pic}^{\mathrm{rel}}_{\Gamma_{d}}\to\mathfrak{Pic}^{\mathrm{rel}}_{g}$ be the boundary stratum corresponding to a decorated prestable graph $\Gamma_{d}$ with two vertices of degree splitting d,$d$,-d, connected by two edges $e_{1}=(h_{1},h_{1}^{\prime})$ and $e_{2}=(h_{2},h_{2}^{\prime})$. Then Pixton’s formula has the form
@@
For our purpose, we further refine the relation $\mathsf{P}_{g}^{g+1}=0$ using the “multiplication by NN” map
$N]:\mathfrak{Pic}_{g}^{\mathrm{rel}}\to\mathfrak{Pic}_{g}^{\mathrm{rel}}.$
@@
$N]^{*}\theta^{g+1}=N^{2g+2}\cdot\theta^{g+1}$
@@
$N]^{*}\left(\theta^{a}i_{*}\big((\psi_{h}+\psi_{h^{\prime}})^{b}\big)\right)=N^{2a}\cdot\theta^{a}i_{*}\big((\psi_{h}+\psi_{h^{\prime}})\big)^{b}$
@@
$N]^{*}\left(\theta^{k}i^{d}_{*}\big((\psi_{h_{1}}+\psi_{h_{1}^{\prime}})^{\ell}(\psi_{h_{2}}+\psi_{h_{2}^{\prime}})^{m}\big)\right)=N^{2k}\cdot\theta^{k}i^{d/N}_{*}\big((\psi_{h_{1}}+\psi_{h_{1}^{\prime}})^{\ell}(\psi_{h_{2}}+\psi_{h_{2}^{\prime}})^{m}\big)$
if $N$ divides $d$ and zero otherwise. Applying $N]^{*}$ to the right-hand side of (33), the first two terms are clearly polynomials in $N$ and for the third term, we have
@@
which is a polynomial in $N$ because $f_{k,\ell,m}(d)$ are polynomials. Therefore we find an expression of $N]^{*}\mathsf{P}^{g+1}_{g}=0$ which is a polynomial in $N$, so that each coefficient gives rise to a relation on $\mathfrak{Pic}^{\mathrm{rel}}_{g}$. We take the coefficient of $N^{2g+2}$ in the expression of $N]^{*}\mathsf{P}_{g}^{g+1}=0$ and pull it back to $\overline{J}_{g}^{\leq 1}$. In particular, the second summand of the right-hand side of (33) does not contribute for weight reasons. This gives the desired relation on $\overline{J}_{g}^{\leq 1}$ of the proposition.
@@ 1. $\bullet$
    genus g1g-1 with degree 1-1, and
@@ 2. $\bullet$
    genus 0 with degree 11
connected by two edges corresponds to a class under the pushforward $\iota_{*}\epsilon_{*}t_{*}$. We denote the two edges by $e_{1}=(h_{1},h_{1}^{\prime})$ and $e_{2}=(h_{2},h_{2}^{\prime})$ where $h_{1},h_{2}$ are the two half-edges attached to the genus g1g-1 vertex. After pulling back to $\overline{J}_{g}^{\leq 1}$, other strata do not contribute by the quasi-stable model of a rank $1$ torsion-free sheaf on a nodal curve. Then we have
@@
We explain how to interpret the classes given by prestable graphs decorated by $\psi$-monomials as classes pushed forward from $J_{g-1,2}$. Let $\mathcal{L}$} be the universal line bundle on the universal curve over $J_{g-1,2}$ trivialized along the first marking, and let $\xi_{2}\in\mathrm{CH}^{1}_{(1)}(J_{g-1,2},{\mathbb{Q}})$ be $c_{1}({\mathcal{L}})$ pulled back along the second marking. Analogous classes are defined on $J_{g-1,2}^{-1}$. We take the coefficient of $N^{1}$ in the expression of $N]^{*}\mathsf{P}^{1}_{0,(1,-1)}=0$. This gives a relation
@@
at the unstable genus 0 vertex. Since $\xi_{h^{\prime}_{1}}=\xi_{h_{1}}=0$ and $\xi_{h^{\prime}_{2}}=\xi_{h_{2}}=\xi_{2}$, the “other terms” in (34) have an expression in terms of monomials of $\theta,\psi_{\bullet},\xi_{2}$. Under the translation, these monomials pull back to classes on $J_{g-1,2}$ of weights <2g2<2g-2. ∎
@@
Although the “multiplication by NN” map extends to the relative Picard stack $\mathfrak{Pic}^{\mathrm{rel}}_{g}$, the cohomology $H^{*}(\mathfrak{Pic}^{\mathrm{rel}}_{g},{\mathbb{Q}})$ may not be a direct sum of weight spaces since it is infinite-dimensional.
We denote by $J^{\leq 1}_{g}\subset\overline{J}^{\leq 1}_{g}$ the open subset parameterizing degree 0 line bundles. We have the tautological relation
$\theta^{g+1}=0\in\mathrm{CH}^{g+1}(J^{\leq 1}_{g},{\mathbb{Q}})$
@@
By Proposition 4.2, it is enough to calculate the pushforward along $\pi$ of the classes $\iota_{*}\epsilon_{*}t_{*}(\theta^{g-1}/(g-1)!)$ and $\iota_{*}\epsilon_{*}t_{*}\alpha$. Since $\alpha\in\mathrm{CH}^{g-1}(J_{g-1,2},{\mathbb{Q}})$ has weights <2g2<2g-2, we have $\pi_{*}\iota_{*}\epsilon_{*}t_{*}\alpha=\iota_{*}\pi^{\prime}_{*}t_{*}\alpha=0$. We also recall
@@
where $0:{\mathcal{M}}_{g-1,2}\to J_{g-1,2}$ is the 0-section; see e.g. [17]. Therefore by the diagram (32) we have
@@
We recall a few standard tautological classes. The cotangent line of the $i$-th marking defines a line bundle ${\mathbb{L}}_{i}$ on $\overline{{\mathcal{M}}}_{g,n}$ with first Chern class
@@
After restriction, we have the $\kappa$-classes on $\overline{\mathcal{M}}_{g}^{\leq 1}$. We also define the boundary class
@@
We consider $J_{g}^{\leq 1}\subset\overline{J}_{g}^{\leq 1}$ the open locus of degree 0 line bundles. Since its complement has (complex) codimension $2$, the restriction $H^{2}(\overline{J}_{g}^{\leq 1},{\mathbb{Q}})\to H^{2}(J_{g}^{\leq 1},{\mathbb{Q}})$ is an isomorphism.
We have the long exact sequence
$\cdots\to H^{0}(J_{g}^{\leq 1}\setminus J_{g},{\mathbb{Q}})\to H^{2}(J_{g}^{\leq 1},{\mathbb{Q}})\to H^{2}(J_{g},{\mathbb{Q}})\to H^{1}(J_{g}^{\leq 1}\setminus J_{g},{\mathbb{Q}})\to\cdots.$
By [18], we know that $H^{2}(J_{g},{\mathbb{Q}})$ is spanned by $\theta$ and $\kappa_{1}$. The lemma follows since the image of $H^{0}(J_{g}^{\leq 1}\setminus J_{g},{\mathbb{Q}})$ is spanned by $\delta$. ∎
@@
Let $\bar{p}:\overline{{\mathcal{C}}}_{g}\to J_{g-1,2}^{-1}$ be the universal curve on $\overline{J}_{g}^{\leq 1}$ pulled back to $J^{-1}_{g-1,2}$ along $\iota\circ\epsilon$, and let $q:\overline{{\mathcal{C}}}^{\mathrm{qs}}_{g}\to\overline{{\mathcal{C}}}_{g}$ be the quasi-stable model. Let ${\mathcal{L}}^{\mathrm{qs}}$ be the universal admissible line bundle on $\overline{{\mathcal{C}}}^{\mathrm{qs}}_{g}$ trivialized along the first marking. The normalization of $\overline{{\mathcal{C}}}^{\mathrm{qs}}_{g}$ is the disjoint union of a ${\mathbb{P}}^{1}$-bundle $\mathbb{P}$} over $J_{g-1,2}$ and the universal curve ${\mathcal{C}}_{g-1,2}$. By [19, eq. (8)], we have a short exact sequence
@@
where we used the self-intersection formula $x_{i}]^{2}=x_{i*}(-\psi_{i})$. ∎
@@
By taking the weight $2g$ part with respect to the “multiplication by NN” map, we get
@@
This shows that $a\kappa_{1}$ is supported on the boundary $\overline{{\mathcal{M}}}_{g}^{\leq 1}\setminus{\mathcal{M}}_{g}$ and is therefore proportional to $\delta$ by the irreducibility of the boundary. [^9] If $g\geq 3$, we can directly deduce $a=0$ from $\kappa_{1}\neq 0\in H^{2}({\mathcal{M}}_{g},{\mathbb{Q}})$.
@@
Since the excess intersection term for $\iota$ is $(\psi_{1}+\psi_{2})$, we have
@@
By taking the weight 2g22g-2 part with respect to the “multiplication by NN” map, we get
@@
Here the coefficient $\frac{1}{48}$ can be uniquely determined by the relation
@@
Therefore, we have to consider the larger open subset $\overline{{\mathcal{M}}}^{\mathrm{int}}_{2}\subset\overline{{\mathcal{M}}}_{2}$ consisting of integral curves and the relative compactified Jacobian $\pi^{\mathrm{int}}:\overline{J}^{\mathrm{int}}_{2}\to\overline{{\mathcal{M}}}^{\mathrm{int}}_{2}$ as in Theorem 4.1(b).
Since the codimension of $\overline{\mathcal{M}}_{2}^{\mathrm{int}}\setminus\overline{{\mathcal{M}}}_{2}^{\leq 1}$ is $2$, Proposition 4.8(a) shows that a generalized theta divisor has to be of the form
@@
$\overline{{\mathcal{M}}}_{1,2}^{\circ}=\overline{{\mathcal{M}}}_{1,2}\setminus\left(\overline{{\mathcal{M}}}_{1,1}\times\overline{{\mathcal{M}}}_{0,3}\right)\subset\overline{{\mathcal{M}}}_{1,2}$
@@
On $\overline{J}^{\mathrm{int}}_{2}$, the universal double ramification cycle relation $\mathsf{P}^{3}_{2}=0$ contains contribution from the boundary strata with two loops. However, those boundary strata contribute to lower degree coefficients of the expression $N]^{*}\mathsf{P}^{3}_{2}=0$. Therefore Proposition 4.2 continues to hold on $\overline{J}^{\mathrm{int}}_{2}$. A direct computation yields $\alpha_{(0)}=\frac{1}{480}(\psi_{1}+\psi_{2})$. Combining with Proposition 4.2 and Lemma 4.7, we obtain
$0$ $=\pi^{\mathrm{int}}_{*}(\theta{\Theta}^{3})=\pi^{\mathrm{int}}_{*}\left(\iota^{\mathrm{int}}_{*}\epsilon_{*}t_{*}\Big(\frac{1}{8}(\theta+\frac{1}{2}(\psi_{1}+\psi_{2}))^{2}+6\theta\alpha_{(0)}-\frac{3b}{8}\theta(\psi_{1}+\psi_{2})\Big)+3b^{2}\theta^{2}\delta^{2}\right)$